what is the sine of the angle whose tangent is the square root of 7?
Answer by Alan3354(69443)

what is the sine of the angle whose tangent is the square root of 7?
-------------
tan = y/x = sqrt(7)/1
r = sqrt(x^2 + y^2) = sqrt(8)
-----
sin = y/r
sin = sqrt(7)/sqrt(8)
= sqrt(56)/8
= sqrt(14)/4
======================
The angle might be in Q1 or Q3, so the sign could be + or minus.